OpenStudy (anonymous):

Yo can I get a little help over here? Find the value of ab when a= 3/4 and b= 1/8

OpenStudy (shinalcantara):

you just have to multiply a and b as stated in the problem \[ab=\frac{ 3 }{ 4 } \times \frac{ 1 }{ 8 } = ?\]

OpenStudy (anonymous):

how do i simplify 7/8 ?

OpenStudy (anonymous):

3 over 12?

OpenStudy (zzr0ck3r):

\(\frac{a}{b}*\frac{c}{d}=\frac{ac}{bd}\)

OpenStudy (zzr0ck3r):

if the greatest common divider of two numbers is 1, then there fraction is in reduced form i.e. notice that only 1 divides both 7 and 8, so \(\frac{7}{8}\) is reduced

OpenStudy (zzr0ck3r):

To reduce 3/12 can we think of a number that divides 3 and 12? hint: it can be one of the numbers.

OpenStudy (anonymous):

3 over 32

OpenStudy (anonymous):

isthat a question?